# Prospective Learning: Learning for a Dynamic Future

### v1

@@ 2.1 Different prospective learning scenarios with illustrative examples
Figure 1: A schematic for prospective learning (left) and realizations of the examples for the four scenarios (top right); dots denote 1s and empty spaces denote 0s for $Y_{t}\in\{0,1\}$ with $X_{t}=1$ for all times $t$. Prospective risk of learners at different times is shown in the bottom panels and discussed in Section 2.1. Scenario 1: For Bernoulli probability $p=0.2$, the maximum-likelihood estimator (MLE) in blue uses a time-agnostic hypothesis $h_{t}(X_{t})=\mathbf{1}(\hat{p}_{t}>0.5)$ where $\hat{p}_{t}=t^{-1}\sum_{s=1}^{t}y_{s}$, ties at $\hat{p}_{t}=0.5$ are broken randomly. The risk of this learner converges to the Bayes risk. Scenario 2: For Bernoulli probability $p=0.2$, the MLE estimator (blue) performs at chance levels. A prospective learner (orange) that alternates between two predictors at even and odd times converges to Bayes risk. Variants of this learner that use less information from the stochastic process (green does not know that the data distributions at even and odd times are tied, red does not know that the distribution shifts at every time-step) also converge to Bayes risk, but more slowly. Scenario 3: For $\theta=0.1$ and $\gamma=0.9$ in the discounted prospective risk, the MLE estimator (blue) again performs at chance levels. A prospective learner that computes an estimate of the transition probability of the two-state Markov chain to estimate $\operatorname{\mathbb{P}}(Y_{t^{\prime}}\mid y_{t})$ for future times $t^{\prime}>t$ converges to Bayes risk. Scenario 4: For $\theta_{0}=\theta_{1}=0.1$, the MLE estimator (blue) performs at chance levels. A prospective learner that uses a variant of Q-learning (described in the text and Section B.3) converges to the prospective Bayes risk.
@@ Acknowledgement
This work was funded by grants from the National Science Foundation (IIS-2145164, CCF-2212519) and the Office of Naval Research (ONR) N00014-22-1-2255. RR was funded by a fellowship from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI. ADS was supported by a fellowship from the Mathematical Institute for Data Science (MINDS) at JHU. We are grateful to all those who provided helpful feedback on earlier drafts of this work, including Marlos Machado.
@@ B.1 Can learning benefit from prospection in the IID scenario?
Figure A.1: Prospective risk of MLE (blue), MAP (purple), and prospective MAP (red) based learners with respect to time. Both MAP and prospective MAP estimators assume a prior distribution of $\text{Beta}(12,16)$ over $p$.
@@ H.1 Prompts for testing prospective learning in LLMs
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide the list of contributions in Section 1 with pointers to this section from the introduction.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Section 7 discusses the limitations of our theoretical and empirical results.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: We have developed substantial new theoretical results in this paper. They are discussed in Section 4. All mathematical claims, illustrative examples, and remarks come with elaborate proofs, both in the main paper and the Appendix.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We use standard visual recognition datasets along with illustrative examples on synthetic tasks. All information to reproduce our experiments, e.g., datasets, network architectures, and training procedures, is provided in Appendix F. All the code that was used to conduct the experiments has been made public.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: All the code that was used to conduct the experiments has been made public.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We have conducted extremely thorough train/test splits, and tuned hyper-parameters manually across multiple runs. This information is provided in Appendix F.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: For all the learning algorithms considered in every experiment in our paper, we compute empirical prospective and instantaneous risks over multiple random seeds, at $15$-$40$ different time steps. This makes our results extremely rigorous and reproducible. All experimental results in the paper come with error bars assuming a Normal distribution for the relevant quantities.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We have conducted experiments on this problem for close to 1.5 years. In addition to the experimental results that are discussed here, there a substantial number of secondary experiments, failed attempts, that are not reported. All the experiments in this paper approximately takes $\sim 200$ GPU hours to run.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We did not use human subjects or participants in these experiments. All our datasets are public ones that are used very commonly in the published literature, or synthetically generated. We do not anticipate societally harmful consequences of this research.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. Justification: This is foundational work and it is not tied to any particular application or deployment. We do not anticipate negative social impact. Down the line, when this line of research matures and it is deployed for making predictions on users, we anticipate a lot of positive social impacts, e.g., prospective learning may enable us to deploy ML models without having to retrain them as data changes over time. This improves performance because it prevents the predictions from deteriorating, but it also leads to other benefits in terms of economic and computational resources necessary to serve ML models.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We only use existing public datasets (appropriately cited in the paper) and synthetic data.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We only use existing public datasets (appropriately cited in the paper) and synthetic data. No new assets are created from this paper.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [N/A]
  52. Justification: No assets are created in this paper.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: No crowdsourcing experiments or research with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: No crowdsourcing experiments or research with human subjects.

### v2

@@ 2.1 Different prospective learning scenarios with illustrative examples
Figure 1: A schematic for prospective learning (left) and realizations of the examples for the four scenarios (top right); dots denote 1s and empty spaces denote 0s for $Y_{t}\in\{0,1\}$ with $X_{t}=1$ for all times $t$. Prospective risk of learners at different times is shown in the bottom panels and discussed in Section 2.1. Scenario 1: For Bernoulli probability $p=0.2$, the maximum-likelihood estimator (MLE) in blue uses a time-agnostic hypothesis $h_{t}(X_{t})=\mathbf{1}(\hat{p}_{t}>0.5)$ where $\hat{p}_{t}=t^{-1}\sum_{s=1}^{t}y_{s}$, ties at $\hat{p}_{t}=0.5$ are broken randomly. The risk of this learner converges to the Bayes risk. Scenario 2: For Bernoulli probability $p=0.2$, the MLE estimator (blue) performs at chance levels. A prospective learner (red) that alternates between two predictors at even and odd times converges to Bayes risk. Variants of this learner that use less information from the stochastic process (purple does not know that the data distributions at even and odd times are tied, green does not know that the distribution shifts at every time-step) also converge to Bayes risk, but more slowly. Scenario 3: For $\theta=0.1$ and $\gamma=0.9$ in the discounted prospective risk, the MLE estimator (blue) again performs at chance levels. A prospective learner that computes an estimate of the transition probability of the two-state Markov chain to estimate $\operatorname{\mathbb{P}}(Y_{t^{\prime}}\mid y_{t})$ for future times $t^{\prime}>t$ converges to Bayes risk. Scenario 4: For $\theta_{0}=\theta_{1}=0.1$, the MLE estimator (blue) performs at chance levels. A prospective learner that uses a variant of Q-learning (described in the text and Section B.3) converges to the prospective Bayes risk.
@@ Acknowledgement
This work was supported by funds provided by the National Science Foundation (IIS-2145164, CCF-2212519), the Office of Naval Research (N00014-22-1-2255), and by the National Science Foundation, DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence), NSF AI Institute Planning award (#2020312), NSF-Simons Research Collaborations on the Mathematical and Scientific Foundations of Deep Learning (MoDL) and THEORINET (#2031985). RR was funded by a fellowship from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI. ADS was supported by a fellowship from the Mathematical Institute for Data Science (MINDS) at JHU. We are grateful to all those who provided helpful feedback on earlier drafts of this work, including Marlos Machado.
@@ B.1 Can learning benefit from prospection in the IID scenario?
Figure A.1: Prospective risk of MLE (blue), MAP (purple), prospective MAP (red) and random-chance (orange) based learners with respect to time. Both MAP and prospective MAP estimators assume a prior distribution of $\text{Beta}(12,16)$ over $p$.